# BPDec: Unveiling the Potential of Masked Language Modeling Decoder in BERT Model Pretraining

### v2

@@ 3.3 Random Mix of Encoder and Decoder Outputs
In our approach, we introduce a modest blending of the encoder’s output into the decoder’s output as shown in Figure 2. This technique serves a dual purpose. Firstly, it encourages the MLM task to not depend exclusively on the abstract representations formed in the deeper layers of the decoder. Instead, it also takes into account the more direct and less processed information from the encoder stage. Drawing inspiration from the NEFTUNE method, which emphasizes the benefits of introducing a certain degree of randomness in Language Model training, we understand that this can potentially enhance performance [14]. Additionally, echoing the Theseus Compression Method utilized in BERT-of-Theseus, the authors recognize that random layer replacement is an effective approach to model compression [11]. In our case, randomly skipping decoder layers can be seen as a form of compression as well. Our experimental results show that while this mixing occurs across batch and sequence dimensions, preserving the structure within the hidden embedding dimension is crucial to maintain the integrity and quality of the output. Experiments also reveal that using 80% of the decoder output mixed with 20% of the encoder output can significantly improve performance on downstream tasks.
@@ 4 Results
In order to examine the benefits of BPDec, we conduct experiments on various NLP tasks with a few models as baselines.
@@ 4.1 Pretraining
Pretraining was conducted on both BERT+BPDec-base and BERT+BPDec-large models. The encoder architecture, pretraining data, pretraining tasks, and mask settings align precisely with those of the original BERT-base and BERT-large models. The distinctive aspect of our BPDec method lies in the addition of a decoder following the encoder. Details about the parameters of this decoder can be found in Table 1. This strategic augmentation aims to enhance the model’s capabilities while maintaining the foundational strengths of BERT, offering a nuanced approach to language modeling that leverages the best of both encoder and decoder functionalities.
Hyperparam BERT+BPDec-base BERT+BPDec-large
Number of decoder layers 2 4
decoder layers with attention mask [1] [1, 2]
decoder layers without attention mask [2] [3, 4]
% of mix with encoder output 20 20
Hidden size 768 1024
FFN inner hidden size 3072 4096
Number of attention heads 12 16
Attention head size 64 64
@@
Table 1: Hyperparameters for decoders used in BERT+BPDec-base and BERT+BPDec-large pretraining. The architectural hyperparameters, including hidden size and the number of attention heads, are consistent with those of the corresponding BERT encoder. These aligned parameters that remain unchanged are highlighted with a gray background in the table for easy reference.
@@
In our study, we conducted a comparative analysis using the BERT model and BERT model with pre-LN [15] trained under identical settings as a baseline, and the DeBERTa model as a point of comparison. Here, we trained our own baseline models, utilizing the same hyperparameters, tokenizer, and pretraining dataset to isolate the impact of other factors on model performance. Our focus was solely on determining the effectiveness of the architectural change. For instance, with RoBERTa [9], authors significantly extended the pretraining duration, increasing the number of steps from 100K to to 500K. This extended training led to substantial improvements in downstream task performance, with the 500K step models outperforming other models in most tasks. Authors emphasized that the longest-trained model still haven’t show signs of overfitting. This indicates that simply increasing training steps can continuously improve the model. Without a fully aligned comparison, we might not arrive at the most insightful conclusions regarding the true impact of architectural modifications on model performance.
@@
Moreover, it is important to note that although the DeBERTa model was trained for the same number of steps as other models, its disentangled attention mechanism introduced a significant overhead comparing with original BERT model. This added complexity affected both the pretraining and fine-tuning phases, but ours BERT+BPDec models only have extra ovearhead during pretraining. The implications of this increased overhead are critical, especially when considering the balance between model performance and computational efficiency. Our analysis aims to provide insights into how these different architectures impact training and operational efficiency, while also maintaining or enhancing language processing capabilities.
@@ 4.2 Performance on GLUE Tasks
BERT+BPDec-base 85.86 83.13 85.66/85.59 88.73 92.39 90.98 71.12 93.81 86.36
@@
In our finetuning experiments, each model was trained on the designated training set and subsequently evaluated on the development set. Table 2 illustrates that under identical conditions and with the same number of pretraining samples, BERT with BPDec enhancement consistently outperforms both BERT and DeBERTa models in terms of the averaged score. Notably, MNLI, the largest dataset among these tasks, serves as a robust indicator of model performance. Our model competes closely with the similarly sized DeBERTa model, achieving slightly higher scores in the matched condition and comparable results in the mismatched one. However, it’s crucial to consider that the DeBERTa model’s encoder is approximately 17% larger in parameter size than BERT encoder, and it incurs a 36% higher cost in finetuning time and a 30% increase in inference time. Despite these disparities, BERT+BPDec matches or even surpasses DeBERTa in various benchmarks in both base and large level sizes. Moreover, BERT+BPDec achieves significant improvement over the original BERT at the same finetuning and serving costs, underlining its efficiency and effectiveness. The results tell us that this decoder side only enhancement in BPDec significantly boosts performance without incurring further additional costs. Building on this, we extended our testing to question answering tasks, further solidifying the viewpoint that our work can yield substantial performance gains.
@@ 4.3 Performance on SQuAD Tasks
DeBERTa-base [12] 83.75 90.78 77.71 80.77
BERT+BPDec-base 84.20 91.12 77.87 80.94
@@
DeBERTa-large [12] 85.91 92.62 82.09 85.12
BERT+BPDec-large 86.11 92.74 82.50 85.29
@@
The experiments conducted on both the v1 and v2 tasks, as presented in Table 3, followed a similar trend to those performed on the GLUE tasks: all models were trained on the training set and evaluated on the development set. While the results for BERT+BPDec on SQuAD v1 did not achieve the highest scores in the table, its performance on SQuAD v2, particularly for the base size model, was notably stronger than the original BERT. Especially, when considering that BERT+BPDec competes closely with, or even surpasses, the DeBERTa model, the improvement is significant. These results underscore BPDec’s efficiency and efficacy, showcasing its ability to deliver robust performance in question answering tasks, a key area in NLP, without additional overhead.
@@ 4.4.1 Number of MLM Decoder Layers
4.4.2 Decoder Layers without Attention Masks
Layers Without Attention Masks AX COLA MNLI-m/mm MRPC QNLI QQP SQuAD v2 EM/F1
[][][ ] 85.10 83.41 85.30/85.06 88.24 91.67 90.78 76.92/80.12
decoder $[1,2]$ 84.77 82.84 84.80/85.08 87.50 91.32 90.86 78.23/81.26
decoder $[2]$ 85.08 83.13 85.28/85.40 89.46 91.62 90.99 78.83/81.73
encoder $[11,12]$ decoder $[1,2]$ 84.95 83.70 84.49/85.02 87.75 91.43 90.86 77.68/80.61
Table 5: Comparative results on various tasks for a 12-layer BERT + 2-layer BPDec model with different configurations regarding the application of attention mask restrictions during pretraining. The finetuning is applied to BERT encoder only. The configurations are denoted by lists, where each list specifies the layers operating without attention mask restrictions.
In our ablation study shown in Table 5, we examined the effects of releasing attention mask restrictions across various configurations of the BPDec model during LM pretraining. Initially, this approach was applied to the encoder layers, where we observed no positive impact on performance. However, a notable improvement emerged when the same restriction release was applied to selected layers of the BPDec decoder. Additional experiments were carried out on the BERT+BPDec-large, revealing that the best results were achieved by removing the attention mask restrictions on the final two layers of a four-layer decoder. Here, we found that the optimal configuration involved releasing the attention mask restrictions in the latter 50% of the decoder layers. This specific adjustment yielded the most significant performance enhancements. These results highlight the nuanced interplay between model architecture and attention mechanisms, emphasizing the potential of targeted modifications in MLM decoder to achieve better outcomes in complex language processing tasks.
@@ 4.4.3 Effectiveness of Output Random Mix
Mix Axis AX COLA MNLI-m/mm MRPC QNLI QQP SQuAD v2 EM/F1
(batch, seq, hidden) 84.68 83.03 84.78/84.91 88.73 91.16 90.80 77.83/80.84
(batch, seq) 85.86 83.13 85.66/85.59 88.73 92.39 90.98 77.87/80.94
(batch) 85.17 83.60 85.11/85.23 88.97 92.46 90.88 78.37/81.34
Table 6: Comparative results on various tasks with different mixing axis configurations on BERT+BPDec-base model during pretraining. We employ the decoder output values 80% of the time, while for the remaining 20%, we utilize the values from the encoder. The configurations specify the dimensions across which mixing is applied, impacting the model’s performance on a range of tasks.
@@
In our ablation studies, we explored the integration of randomness into the model output before the Softmax layer with Masked Language Modeling Task. A particularly effective method was found is to randomly alternate between using outputs from the decoder and the encoder. Some results of these experiments are detailed in Table 6. We observed that a complete mix of the two outputs did not yield optimal results, possibly due to the disruption of the output embeddings. To address this, we adjusted our randomization strategy to only apply randomness along the (batch, seq) dimensions, maintaining the cohesiveness of individual output embeddings. This implies that an output embedding is sourced entirely from either the decoder or the encoder.
Furthermore, we experimented with varying mix ratios of decoder and encoder outputs. As shown in Table 7, the most effective ratio during pretraining was found to be 80% decoder output and 20% encoder output. This specific mix ratio led to the best performance, underlining the importance of fine-tuning the balance between the two types of outputs to achieve optimal results in the finetuning phase. The 80% ratio for utilizing decoder output emerged as a reasonable choice. If this ratio were lower, it might result in insufficient training of the decoder, diminishing its potential benefits. Conversely, if the ratio were higher, it could undermine the intended effect of randomness. Therefore, this balanced approach not only ensures effective training of the decoder but also maintains the necessary level of randomness to optimize the model’s overall performance.
@@ 5 Conclusion
In this paper, we introduced BERT+BPDec, an enhanced version of the original BERT model, focusing on improvements with the Masked Language Modeling (MLM) Decoder used in pretraining. Key innovations include adding transformer block layers as MLM decoder, removing restrictions on attending to masked positions in the MLM decoder, and introducing a degree of output randomness by combining encoder and decoder outputs. Through rigorous evaluations on tasks such as MNLI, SQuAD, and RACE, BERT+BPDec model demonstrated superior performance over the original BERT and other state-of-the-art models, without any increasing computational complexity in further finetuning and inference. The ablation study confirmed the effectiveness of each modification, highlighting BPDec’s potential in advancing NLP efficiency and effectiveness. Furthermore, we provided optimal hyperparameter settings for BPDec, aiding in enhancing efficiency, ensuring consistent performance, while also reducing energy consumption and emissions. This work not only presents a significant step forward in language model optimization but also offers valuable insights for future research in the field.

### v3

@@ 3.3 Random Mix of Encoder and Decoder Outputs
In our approach, we introduce a modest blending of the encoder’s output into the decoder’s output as shown in Figure 3. This technique serves a dual purpose. Firstly, it encourages the MLM task to not depend exclusively on the representations formed in the deeper layers of the decoder. Instead, it also takes into account the more direct and less processed information from the encoder stage. Drawing inspiration from the NEFTUNE method, which highlights the potential performance benefits of introducing controlled randomness during language model training, we adopt a similar principle in our approach [15]. Additionally, echoing the compression method utilized in BERT-of-Theseus, the authors recognize that random layer replacement is an effective approach to model compression [12]. In our case, randomly skipping decoder layers can be seen as a form of compression as well. Our experimental results reveal that using 80% of the decoder output mixed with 20% of the encoder output can significantly improve performance on downstream tasks.
@@ 4 Results
In order to examine the benefits of BPDec, we conduct experiments on various NLP tasks with a few models with as baselines.
@@ 4.1 Pretraining
Pretraining was conducted on both base and large model sizes. Additionally, we introduced a variant of the base model (base-h256) to further demonstrate the generalizability of our method on smaller models. The encoder architecture, pretraining data, pretraining tasks, and mask settings are identical with the original BERT models. Our BPDec lies in the addition of a decoder following the encoder. Details about the parameters of this decoder can be found in Table 1. This strategic augmentation aims to enhance the model’s capabilities while maintaining the foundational architecture of BERT encoder, and it also ensures a high degree of fairness in our comparative experiments.
BERT+BPDec Hyperparam base-h256 base large
Number of decoder layers 2 2 4
decoder layers with unmasking [1, 2] [1, 2] [1, 3]
decoder layer unmasking rates [0.5, 1.0] [0.5, 1.0] [0.5, 1.0]
% of mix with encoder output 20 20 20
Number of encoder layers 12 12 24
Hidden size 256 768 1024
FFN inner hidden size 1024 3072 4096
Number of attention heads 4 12 16
Attention head size 64 64 64
@@
Table 1: Hyperparameters for decoders used in BERT+BPDec-base-h256, BERT+BPDec-base and BERT+BPDec-large pretraining. The architectural hyperparameters, including hidden size and the number of attention heads, are the exaclty same with the corresponding BERT model. These parameters that remain unchanged are highlighted with a gray background in the table for easy reference.
@@
In our study, we conducted a comparative analysis using the BERT model, BERT model with pre-LN (pre-layer normalization) [16], and DeBERTa [13] trained under identical settings as baselines. Here, we utilized the same hyperparameters, tokenizer, and pretraining dataset to isolate the impact of other factors on model performance. Our focus was solely on determining the effectiveness of the architectural change. In RoBERTa [9], authors significantly extended the pretraining duration, increasing the number of steps from 100K to 500K. This led to substantial improvements in downstream task performance. Authors also emphasized that the longest-trained model still haven’t show signs of overfitting. This observation highlights the importance of controlling for training steps and pretraining dataset when evaluating architectural modifications, as it can act as a confounding factor that influences model performance. Without careful consideration of such factors, drawing definitive conclusions about the true impact of architectural changes could be misleading.
@@
Moreover, it is important to note that although the DeBERTa model was trained for the same number of steps as other models, its disentangled attention mechanism introduced a significant overhead comparing with other baseline models and with our BERT+BPDec models. This added complexity affected both the pretraining and fine-tuning phases, but ours BERT+BPDec models only have extra ovearhead during pretraining. The implications of this increased overhead are critical, especially when considering the balance between model performance and computational efficiency. Our analysis aims to provide insights into how these different architectures impact training and computation efficiency, while also maintaining or enhancing language processing capabilities.
@@ 4.2 Performance on GLUE Tasks
BERT-base-h256 [1] 78.99 75.26 79.09/79.83 81.86 87.81 88.45 61.37 87.73 80.04
BERT-PreLN-base-h256 [16] 76.70 69.13 76.48/77.87 71.81 83.40 86.59 55.96 86.01 75.99
DeBERTa-base-h256 [13] 79.04 74.50 79.41/79.99 78.19 87.88 88.68 64.62 88.76 80.11
BERT+BPDec-base-h256 79.02 75.74 79.08/80.33 81.86 88.05 88.46 62.09 89.33 80.44
@@
BERT+BPDec-base 85.76 83.13 85.66/85.59 88.73 92.39 90.98 71.12 93.81 86.35
@@
In our finetuning experiments, each model was trained on the designated training set and subsequently evaluated on the development set. Table 2 illustrates that under identical conditions and with the same number of pretraining samples, BERT with BPDec enhancement consistently outperforms both BERT and DeBERTa models in terms of many tasks and the averaged score. Notably, MNLI, the largest dataset among these tasks, serves as a robust indicator of model performance. Our model competes closely with the similarly sized DeBERTa model, achieving slightly higher scores in the matched condition and comparable results in the mismatched one. However, it’s crucial to consider that the DeBERTa model’s encoder is approximately 17% larger in parameter size than BERT encoder, and it incurs a 32% higher cost in finetuning cost and a 29% increase in inference latency. Despite these disparities, BERT+BPDec matches or even surpasses DeBERTa in various benchmarks in base-h256, base and large level sizes. Moreover, BERT+BPDec achieves significant improvement over the original BERT at exactly the same finetuning and serving costs. The results tell us that this decoder only enhancement significantly boosts performance without incurring further additional costs. Building on this, we extended our testing to question answering tasks, further solidifying the conclusion that our work can yield substantial performance gains.
@@ 4.3 Performance on SQuAD Tasks
BERT-base-h256 [1] 75.05 83.55 67.07 69.75
BERT-PreLN-base-h256 [16] 74.18 80.26 65.39 67.96
DeBERTa-base-h256 [13] 75.91 84.31 67.32 69.87
BERT+BPDec-base-h256 75.74 84.14 67.97 70.82
@@
DeBERTa-base [13] 83.75 90.78 77.87 80.94
BERT+BPDec-base 84.20 91.12 78.11 81.23
@@
DeBERTa-large [13] 86.91 92.96 82.09 85.12
BERT+BPDec-large 86.11 92.70 82.12 85.39
@@
The experiments conducted on both the v1 and v2 tasks, as presented in Table 3, followed a similar trend to those performed on the GLUE tasks: all models were trained on the training set and evaluated on the development set. While the results for BERT+BPDec on SQuAD v1 did not achieve the highest scores in the table, its performance on SQuAD v2, particularly for the base and base-h256 size model, was notably stronger than the original BERT. Especially, when considering that BERT+BPDec competes closely with, or even surpasses, the DeBERTa model, the improvement is significant. These results underscore BPDec’s efficiency and efficacy, showcasing its ability to deliver robust performance in question answering tasks, a key area in NLP, without additional overhead.
@@ 4.4.1 Number of MLM Decoder Layers
4.4.2 Decoder Layers with Attention Mask Relaxation
Decoder layer unmasking rates AX COLA MNLI-m/mm MRPC QNLI QQP SQuAD v2 EM/F1
$[0.0,0.0]$ 85.10 83.41 84.82/85.53 88.97 91.31 90.99 77.67/81.13
$[1.0,N/A]$ 84.77 82.84 84.80/85.08 87.50 91.32 90.86 77.68/80.82
$[0.0,1.0]$ 84.95 83.70 84.49/85.02 87.75 91.43 90.86 78.23/81.26
$[0.5,1.0]$ 85.08 83.13 85.08/85.40 88.46 91.62 90.99 78.83/81.73
Table 5: Comparative results on various tasks for a 12-layer BERT + 2-layer decoder model with different configurations regarding the application of attention mask restrictions during pretraining. The finetuning is applied to BERT encoder only. The configurations are denoted by lists, where each list specifies 2 random unmasking rates for the first and second decoder layers
In our ablation study shown in Table 5, we examined the effects of releasing attention mask restrictions across various configurations of the BPDec model during the MLM pretraining. Initially, we removed the attention mask entirely for the decoder, where we observed no positive impact on performance. However, a notable improvement emerged when the same restriction release was applied to a deeper layer of the BPDec decoder. Moreover, we found an optimal configuration when we gradually remove the mask, the first layer randomly unmask 50%, and the second decoder layer release the 50% attention mask left. This specific adjustment yielded the most significant performance enhancements. These results highlight the nuanced interplay between model architecture and attention mechanisms, emphasizing the potential of targeted modifications in MLM decoder to achieve better outcomes in complex language processing tasks. Additionally, we observed that the improvement on the SQuAD task after adding attention mask relaxation was significantly larger than the gains on other classification tasks. Considering that SQuAD tasks generally have longer and more complex contexts, this indirectly indicates that our improvement can enhance the model’s ability to understand context.
@@ 4.4.3 Effectiveness of Output Random Mix
In our ablation studies, we explored the integration of randomness into the model output before the Softmax layer with Masked Language Modeling Task. A particularly effective method was found is to randomly alternate between using outputs from the decoder and the encoder. We set our randomization strategy to only apply randomness along the (batch, seq) dimensions, maintaining the cohesiveness of individual output embeddings. This implies that an output embedding is sourced entirely from either the decoder or the encoder.
Furthermore, we experimented with varying mix ratios of decoder and encoder outputs. As shown in Table 6, the most effective ratio during pretraining was found to be 80% decoder output and 20% encoder output. This specific mix ratio led to the most balanced performance, underlining the importance of fine-tuning the balance between the two types of outputs to achieve optimal results in the finetuning phase. The 80% ratio for utilizing decoder output emerged as a reasonable choice. If this ratio were lower, it might result in insufficient training of the decoder, diminishing its potential benefits. Conversely, if the ratio were higher, it could undermine the intended effect of randomness.
From the perspective of experimental results, although completely avoiding random output mix yielded the best performance on the SQuAD task, random output mix provides us with an excellent method to balance the model’s capabilities across different tasks. We believe that the 80% rate trades off a portion of SQuAD’s context understanding ability in exchange for better performance on other GLUE tasks, ultimately achieving a good balance. Therefore, this balanced approach not only ensures effective training of the decoder but also maintains the necessary level of randomness to optimize the model’s overall performance.
4.5 Generalizability to Other BERT-like Models
To further demonstrate the effectiveness of BPDec, we applied the same methodology to ALBERT [17] and BERT-PreLN [16] models. For ALBERT, the BPDec does not share parameters with the ALBERT encoder, but the decoder layers within BPDec share parameters. For BERT-preLN, we followed the encoder’s structure and used pre-LayerNorm in BPDec. Experiments on the GLUE benchmark and SQuAD dataset showed consistent performance improvements across both models.
Model AX COLA MNLI-m/mm MRPC QNLI QQP SQuAD v2 EM/F1
ALBERT-base 84.34 80.15 84.46/84.51 85.78 91.03 90.81 75.62/78.77
ALBERT+BPDec-base 84.43 81.78 84.82/85.09 87.75 92.04 90.90 76.71/79.93
BERT-PreLN-base 85.03 83.89 85.10/85.45 88.24 91.96 90.78 77.03/79.87
BERT-PreLN+BPDec-base 85.42 85.08 85.55/85.86 88.75 91.71 90.77 78.04/80.92
Table 7: Comparative results of ALBERT and BERT-PreLN models with and without BPDec.
These additional experiments shown in 7 prove the effectiveness, adaptability, and generalizability of the BPDec method, highlighting its potential for broader application in various BERT-like architectures. The consistent improvements observed across different model types suggest that BPDec’s principles can be successfully transferred to enhance the performance of a wider range of models.
@@ 5 Conclusion
In this paper, we introduced BERT+BPDec, an enhanced version of the original BERT model, focusing on improvements with the Masked Language Modeling (MLM) Decoder used in pretraining. Key innovations include adding transformer block layers as MLM decoder, removing restrictions on attending to masked positions in the MLM decoder, and introducing a degree of output randomness by randomly mixing encoder and decoder outputs. Through rigorous evaluations on tasks such as MNLI, SQuAD, and RACE, BERT+BPDec model demonstrated superior performance over the original BERT and other state-of-the-art models, without any increasing computational complexity in further finetuning and inference. The ablation study confirmed the effectiveness of each modification, highlighting BPDec’s potential in advancing NLP efficiency and effectiveness. The optimal hyperparameter settings we provide for BPDec not only improve the model’s performance and efficiency but also contribute to sustainable computing by reducing energy consumption and emissions. This work has practical implications for real-world applications of language models and paves new avenues for future research on optimizing the masked language modeling pretraining process.